# Plasticine: A Traceable Diffusion Model for Medical Image Translation

### v1

@@ 4.3 Datasets and Evaluation Metrics
Evaluation Metrics To evaluate segmentation performances, we used three metrics: intersection-over-union (mIoUmIoU) [56], the mean Sørensen–Dice coefficient (DiceDice), the accuracy (AccAcc). To evaluate translation performances, we measured the intensity distribution between synthesised and target images by using three metrics: Fréchet inception distance (FIDFID) [57], Simplified Fréchet Distance (SFDSFD) [58] and Bhattacharyya distance $D_{Bhat}$ [59].
@@ 4.4 Traceability Evaluation with Segmentation
Results on segmentation. TABLE 1 show the quantitative segmentation results of compared methods. Among the compared methods, the proposed Plasticine outperforms the previous state-of-the-art methods, i.e., GAN based methods and diffusion based method (DDIB and UNSB), and obtains the highest AccAcc, DiceDice, and mIoUmIoU on retinal OCT and chest MRI-CT datasets. When compared on the cardiac MRI dataset, though the AccAcc of Plasticine is 0.001 lower than that of CycleGAN, Plasticine outperforms all the compared methods in terms of DiceDice with clear improvements.
@@ 4.5 Comparison on Image Synthesis
Results on image synthesis. TABLE 3 shows the quantitative results of image synthesis based on image distribution measurements. In the retinal OCT dataset, our Plasticine outperforms all compared models in all metrics and gains 23.49, 28.67 and 0.012 difference in FIDFID, SFDSFD and $D_{Bhat}$ than that of MUNIT, respectively. For the chest MRI-CT dataset, Plasticine obtains the best FIDFID score and is close to MUNIT on SFDSFD and $D_{Bhat}$ with a difference of 3.91 and 0.008. The visual results shown in Fig. 10 indicate that the images transferred by Plasticine for the chest and cardiac datasets are comparable in shape and intensity translation with that of MUNIT. Although MUNIT can generate images that are close to the target distribution, the generated images are not consistent with the original source image. Additionally, the misaligned segmentation results are not satisfying as shown in Fig. 9. Specifically, the myocardium and ventricles of the heart are not accurately segmented in their proper positions by MUNIT. In the chest dataset, results obtained by MUNIT show a mismatch in the size of the labelled liver and heart.

### v2

@@ 4.3 Datasets and Evaluation Metrics
Evaluation Metrics To evaluate segmentation performances, we used three metrics: intersection-over-union ($mIoU$) [56], the mean Sørensen–Dice coefficient ($Dice$), the accuracy ($Acc$). To evaluate translation performances, we measured the intensity distribution between synthesised and target images by using three metrics: Fréchet inception distance ($FID$) [57], Simplified Fréchet Distance ($SFD$) [58] and Bhattacharyya distance $D_{Bhat}$ [59].
@@ 4.4 Traceability Evaluation with Segmentation
Results on segmentation. TABLE 1 show the quantitative segmentation results of compared methods. Among the compared methods, the proposed Plasticine outperforms the previous state-of-the-art methods, i.e., GAN based methods and diffusion based method (DDIB and UNSB), and obtains the highest $Acc$, DiceDice, and $mIoU$ on retinal OCT and chest MRI-CT datasets. When compared on the cardiac MRI dataset, though the $Acc$ of Plasticine is 0.001 lower than that of CycleGAN, Plasticine outperforms all the compared methods in terms of $Dice$ with clear improvements.
@@ 4.5 Comparison on Image Synthesis
Results on image synthesis. TABLE 3 shows the quantitative results of image synthesis based on image distribution measurements. In the retinal OCT dataset, our Plasticine outperforms all compared models in all metrics and gains 23.49, 28.67 and 0.012 difference in $FID$, SFDSFD and $D_{Bhat}$ than that of MUNIT, respectively. For the chest MRI-CT dataset, Plasticine obtains the best $FID$ score and is close to MUNIT on $SFD$ and $D_{Bhat}$ with a difference of 3.91 and 0.008. The visual results shown in Fig. 10 indicate that the images transferred by Plasticine for the chest and cardiac datasets are comparable in shape and intensity translation with that of MUNIT. Although MUNIT can generate images that are close to the target distribution, the generated images are not consistent with the original source image. Additionally, the misaligned segmentation results are not satisfying as shown in Fig. 9. Specifically, the myocardium and ventricles of the heart are not accurately segmented in their proper positions by MUNIT. In the chest dataset, results obtained by MUNIT show a mismatch in the size of the labelled liver and heart.